You have $15,000 to invest and would like to create a portfolio with an expected return of 9.9 percent. You can invest in Stock K with an expected return of 8.4 percent and Stock L with an expected return of 12 percent. How much will you invest in Stock K?

  • $6,250.00

  • $8,020.83

  • $8,333.33

  • $8,750.00

  • $5,468.75
Answer #1

Let the investment in stock K be x. The total portfolio investment is $15,000. Therefore, the investment in stock L will be ($15,000-X).

The total expected return from portfolio is 9.9% whereas expected returns of stock K and L are 8.4% and 12% respectively. Using the formula of total expected return of a portfolio, we get as follows:

Total expected return= [( Expected return of stock K * weight of stock K in portfolio) + (Expected return of stock L*weight of stock L in portfolio)]/(sum of weights of stock K and stock L)

Replacing the values, we get as follows:-

9.9%= [(8.4%*x) + (12%*(15000-x))]/(x+15000-x)

0.099*15000= 0.084*x + 0.12*15000 - 0.12*x

1485= 0.084x + 1800 - 0.12x

0.12x-0.084x= 1800-1485

0.036x= 315

x= 8750

Therefore, investment in stock K should be $8,750 and the correct option is option D.